OpenStudy (anonymous):

3x(squared)+6x-15=0 solve by completing the square??

OpenStudy (zzr0ck3r):

3x^2 + 6x-15=0 devide away the 3 x^2 +2x -5 = 0 move five over x^2+2x = 5 take half the b term and square it and add to both sides and use that number do make the binomial (x+1)^2 = 6 x+1 = +- sqrt(6) x = sqrt(6) -1 or -sqrt(6) - 1

OpenStudy (zzr0ck3r):

when I say use that number for the binomial i mean 1/2 the b term not 1/2 the b term squared. so if it was like x^2 + 4x = 7 you would do (x+2)^2= 11. the first 2 is half the b term and the 11 is 7 + half the b term squared.